Consider the reaction when aqueous solutions of magnesium chloride and sodium phosphate are combined. The net ionic equation for this reaction is:
1
Expert's answer
2019-03-05T06:13:58-0500

3MgCl2 + 2Na3PO4 = Mg3(PO4)2↓ + 6NaCl

3Mg2+ + 6Cl- + 6Na+ + 2PO43- = Mg3(PO4)2↓ + 6Na+ + 6Cl-

3Mg2+ + 2PO43- = Mg3(PO4)2